A wave is modeled with the function y(x,t) = (0.24m) cos((0.36rad/m)x − (0.71rad/s)t + π/3). Find the wavelength.

Round your answer to 2 decimal places


1
Expert's answer
2021-04-28T13:04:34-0400
"\lambda=\frac{2\pi}{k}\\\lambda=\frac{2\pi}{0.36}=17\ m"